# June 25, 2025

### v1

@@
The LHCb detector underwent a comprehensive upgrade in preparation for the third data-taking run of the Large Hadron Collider (LHC), known as LHCb Upgrade I. The increased data rate of Run 3 not only posed data collection (Online) challenges but also significant Offline data processing and analysis ones. The offline processing and analysis model was consequently upgraded to handle the factor 30 increase in data volume and the associated demands of ever-growing analyst-level datasets, led by the LHCb Data Processing and Analysis (DPA) project. This paper documents the LHCb “Sprucing” — the centralised offline processing, selections and streaming of data — and “Analysis Productions” — the centralised and highly automated declarative nTuple production system. The DaVinci application used by analysis productions for tupling spruced data is described as well as the apd and lbconda tools for data retrieval and analysis environment configuration. These tools allow for greatly improved analyst workflows and analysis preservation. Finally, the approach to data processing and analysis in the High-Luminosity Large Hadron Collider (HL-LHC) era — LHCb Upgrade II — is discussed.
@@ 1 Introduction
  • Sprucing – Centralised offline data skimming, slimming and streaming to reduce the data footprint between tape and disk storage.
  • Analysis productions – Centralised nTuples production using the LHCbDirac [4] transformation system with maximal automation and optimised user experience.
@@ 2.1 The Role of the Sprucing in LHCb Upgrade I
Finally, the TURCAL stream contains selection lines for physics channels that are used for, among other studies, detector calibration. Hence these lines additionally persist raw detector information (CALO) or MUON from instance). In the case of TURCAL, a factor 8 reduction is again required by the Sprucing and the data is further streamed. The TURCAL Sprucing achieves this by facilitating a line-by-line customisation of the persistency of reconstruction objects and detector RawBanks to reduce the data footprint to disk. The TURCAL stream, like the FULL stream, maintains the option for re-processings and even re-running the reconstruction due to the presence of raw detector information. At the end of the 2024 data taking period, the TURCAL stream was running 147 lines, which were subsequently Spruced.
@@ 2.2 Sprucing Productions and Campaigns
Similarly to Stripping of Runs 1&2 [16], the Sprucing runs concurrently with data taking. Multiple Sprucing campaigns are established over a data-taking year corresponding to changes in the configuration of HLT{1,2} or the Sprucing itself that impacts the physics output, for example, updates to existing selection lines or the addition of new ones. The development schedule of trigger and Sprucing campaigns is closely aligned and coordinated.
@@
In ongoing campaigns, the observed turnaround between the HLT2 processing, file transfer to the Offline system, and the Sprucing processing is 24-48 hours.
@@
Re-Sprucing campaigns take place in end-of-year LHC shutdowns when the Sprucing input data can be staged from tape to disk and Offline resources have capacity. This (typically) four-month period constitutes a strict time frame within which the re-Sprucing must be performed to avoid conflict with concurrent campaigns.
Given the full event reconstruction persisted in the FULL stream, re-Sprucings offer the opportunity to re-optimise selections or add new selection lines based on the needs of physics analysts. The preparation of re-Sprucing campaigns starts several weeks before the end of data-taking of the year, when analysts submit requests and implement changes to selection lines. Once the concurrent data processing has finished, the Sprucing input data is staged from tape to disk. A small fraction of data is re-Spruced as part of the campaign validation with Analysis Productions, again, used by analysts to check distributions of variables. Once the green-light is given, the full production is launched in order to complete the processing before the start of the next year’s data-taking.
Re-Sprucing campaigns can either be full — whereby all Sprucing lines are rerun and the concurrent campaign can be superseded — or incremental — where only new or updated lines run and this dataset is additional to the previous full (re-)Sprucing dataset.
@@ 2.3 File Summary Records (FSRs)
A File Summary Record (FSR) is a per-file tree data structure containing information on the file content, stored within the file itself. FSRs cannot be used in the LHCb MDF file format due to their “per-event” storage as described in Section 2.1, so Sprucing is the first point in the data flow in which FSRs can be created. FSRs have been used since LHCb Run 1 to record luminosity information. The total number of luminosity events in a file, identified by a unique routing bit, is recorded inside the FSR to indicate the LHCb luminosity a file represents. Storing this information in an FSR means that these events can be discarded at the Sprucing stage.
In Run 3 the FSR information was expanded to include the decoding tables and the application options of the last processing step making Spruced files fully self-contained. The information is written in JSON format. As described above, LHCb data contains a “DstData” RawBank for physics candidate information and a “HltDecReports” RawBank for the trigger decisions of each event. To save disk space, this data is encoded. The (potentially long) strings of physics object locations (PackedObjectLocations) and the selection line decision names (HLT1SelectionID, HLT2SelectionID, SpruceSelectionID) are mapped to integer values. In the RawBanks, only the integers and corresponding information are stored per event. Thus, in order to read the data, these integer-to-string mappings, referred to as decoding tables, are necessary. Each decoding table is identified by a unique hexadecimal decoding key and stored. To make the Sprucing output files self-contained, the decoding tables are now written to the FSR and can be read from there in further processing steps.
The application options used to create a Spruced file are also now stored in the file’s FSR — this is the Moore options used to run the Sprucing job. Besides the provenance aspect of knowing exactly how a file was produced via metadata stored in the file itself, storing these options allows further automation of Analysis Productions (described in Section 3), as well as configuration consistency checks, as configuration options to run the DaVinci application on Spruced files can be deduced from the FSR. These configuration options include detector geometry and conditions tags, the file type, the input process (e.g., HLT2) and whether the file contains real or simulated data.
@@ 3.3 Production Testing and Submission
Should any monitored criteria exceed predefined thresholds, tests will fail, and feedback is provided to the user (merge request) on the AP GitLab platform. The productions, their status, and details of the validation are summarised on the platform, connected to aforementioned services, with curated job logs highlighting potentially occurring warnings and errors.
@@ 3.5 The apd Tool and Analysis Preservation
The apd tool provides a simple programming interface to the Analysis Productions system. It is a Python package published to standard repositories (such as PyPI [23] and conda-forge [24] that allows the look-up of the physical file names of files created in the context of an Analysis Production, filtering them according to associated metadata (tags). Some of these tags are automatically created by the system (e.g., the data type), but analysts can also define custom tags that indicate the dataset properties or its intended use. Using apd, analysts can avoid keeping lists of physical file names that have time-limited validity and instead declare input data with information that represents its properties and intended use (e.g., the analysis name, the working group it is attached to, the data-taking year, the event type in case of simulated data, etc.).
@@
The ease of use is crucial to help with the adoption of any new system. The apd tool is installed in the default LHCb analysis environment available through lb-conda, the LHCb Conda wrapper script tool [26] that provides access to Conda versions on CERN’s cvmfs [27]. The apd tool has few dependencies and can be installed on any machine independently of the other LHCb software applications if necessary. Using apd from Python scripts requires minimal effort from analysts, and also provides the functionality to cache metadata and files locally, to improve performance, or reproduce part of the analysis locally.
@@
Combining the use of apd to search for LHCb Analysis Production data files, lb-conda to ensure the reproducibility of the analysis environments and Snakemake to define the analysis workflows ensures comprehensive and reliable analysis preservation. The ease of use of this system promoted significant uptake from LHCb analysts.
@@ 4 The DaVinci Application
Analysis Productions uses the offline analysis software project DaVinci to create nTuples for further high-level analysis. The DaVinci application is built on the Gaudi [32] framework software package for processing High Energy Physics event data. Run 3 DaVinci employs the purpose-built FunTuple component to facilitate the storage of event data in ROOT format, optimising it for subsequent offline analysis [33]. These nTuples may contain information about the selected signal candidates and the decay products (particle kinematics, particle identification hypotheses, vertex fit qualities, etc.). Within DaVinci one can run the DecayTreeFitter algorithm [34] on the signal candidates to constrain the particle momenta to point to particular production vertices and constrain the composite particle masses to improve the resolution of the reconstructed objects.
@@ 5 LHCb Offline Data Processing and Analysis in LHCb Upgrade II
  • Binning of variables.
Generally, these steps are trivially parallelisable and are well suited to grid productions. WLCG resources are experiment agnostic and can run any code/tool that can be packaged to cvmfs or deployed via a container. Portable analysis environments thus allow APs to run any application required for the above steps. In particular, the LHCb lb-conda tool that provides access to conda environments in cvmfs offers a comprehensive (versioned) default environment with all common HEP software packages, as well as the LHCb calibration tools. Alongside this, analysts can create customised environments specifically for their analysis, which are versioned, enabling full analysis environment preservation.

### v2

# June 30, 2025

@@
The LHCb detector underwent a comprehensive upgrade in preparation for the third data-taking run of the Large Hadron Collider (LHC), known as LHCb Upgrade I. With its increased data rate, Run 3 introduced considerable challenges in both data acquisition (online) and data processing and analysis (offline). The offline processing and analysis model was upgraded to handle the factor 30 increase in data volume and the associated demands of ever-growing datasets for analysis, led by the LHCb Data Processing and Analysis (DPA) project. This paper documents the LHCb “Sprucing” — the centralised offline processing, selections and streaming of data — and “Analysis Productions” — the centralised and highly automated declarative nTuple production system. The DaVinci application used by analysis productions for tupling spruced data is described as well as the apd and lbconda tools for data retrieval and analysis environment configuration. These tools allow for greatly improved analyst workflows and analysis preservation. Finally, the approach to data processing and analysis in the High-Luminosity Large Hadron Collider (HL-LHC) era — LHCb Upgrade II — is discussed.
@@ 1 Introduction
Sprucing: Centralised offline data skimming, slimming and streaming to reduce the data footprint between tape and disk storage.
Analysis productions: Centralised nTuples production using the LHCbDirac [4] transformation system with maximal automation and optimised user experience.
@@ 2.1 The Role of the Sprucing in LHCb Upgrade I
Finally, the TURCAL stream contains selection lines for physics channels that are used for, among other studies, detector calibration. Hence these lines additionally persist raw detector information (CALO or MUON for instance). In the case of TURCAL, a factor 8 reduction is again required by the Sprucing and the data is further streamed. The TURCAL Sprucing achieves this by facilitating a line-by-line customisation of the persistency of reconstruction objects and detector RawBanks to reduce the data footprint to disk. The TURCAL stream, like the FULL stream, maintains the option for re-processings and even re-running the reconstruction due to the presence of raw detector information. At the end of the 2024 data taking period, the TURCAL stream was running 147 lines, which were subsequently Spruced.
@@ 2.2 Sprucing Productions and Campaigns
Similarly to Stripping of Runs 1 & 2 [16], the Sprucing runs concurrently with data taking. Multiple Sprucing campaigns are established over a data-taking year corresponding to changes in the configuration of HLT1, HLT2 or the Sprucing itself that impacts the physics output, for example, updates to existing selection lines or the addition of new ones. The development schedule of trigger and Sprucing campaigns is closely aligned and coordinated.
@@
In nominal, concurrent campaigns, the observed turnaround between the HLT2 processing, file transfer to the Offline system, and the Sprucing processing is 24-48 hours.
@@
Re-Sprucing campaigns take place in end-of-year LHC shutdowns when the Sprucing input data can be staged from tape to disk and Offline resources have capacity. This (typically) four-month period constitutes a strict time frame within which the re-Sprucing must be completed to avoid conflict with concurrent campaigns.
Given the full event reconstruction persisted in the FULL stream, re-Sprucings offer the opportunity to re-optimise selections or add new selection lines based on the needs of physics analysts. The preparation of re-Sprucing campaigns starts several weeks before the end of data-taking of the year, when analysts submit requests and implement changes to selection lines.
Once the concurrent data processing has finished, the Sprucing input data is staged from tape to disk. A small fraction of data is re-Spruced as part of the campaign validation with Analysis Productions, again, used by analysts to check distributions of variables. Once the green-light is given, the full production is launched in order to complete the processing before the start of the next year’s data-taking.
Re-Sprucing campaigns can either be full — whereby all Sprucing lines are rerun and the concurrent campaign can be superseded — or incremental — where only new or updated lines run and this dataset is additional to the previous full (re-)Sprucing dataset. The decision between running an incremental or full campaign is made based on the number of lines added/modified and the resulting storage requirements.
@@ 2.3 File Summary Records (FSRs)
A File Summary Record (FSR) is a per-file tree data structure containing information on the DST file content, stored within the DST file itself. FSRs cannot be used in the LHCb MDF file format due to their “per-event” storage as described in Section 2.1, so Sprucing is the first point in the data flow in which FSRs can be created. FSRs have been used since LHCb Run 1 to record luminosity information. The total number of luminosity events in a file, identified by a unique routing bit, is recorded inside the FSR to indicate the LHCb luminosity a file represents. Storing this information in an FSR means that these events can be discarded at the Sprucing stage.
In Run 3 the FSR information was expanded to include the decoding tables and the application options of the last processing step making Spruced files fully self-contained. As described above, LHCb data contains a “DstData” RawBank for physics candidate information and a “HltDecReports” RawBank for the trigger decisions of each event. To save disk space, this data is encoded. The (potentially long) strings of physics object locations (PackedObjectLocations) and the selection line decision names (HLT1SelectionID, HLT2SelectionID, SpruceSelectionID) are mapped to integer values. In the RawBanks, only the integers and corresponding information are stored per event and so, in order to read the data, these integer-to-string mappings, referred to as decoding tables, are necessary. Each decoding table is identified by a unique hexadecimal decoding key and stored. To make the Sprucing output files self-contained, the decoding tables are now written to the FSR and can be read from there in further processing steps.
The application options used to create a Spruced file are also now stored in the file’s FSR — these are the Moore options used to run the Sprucing job. Besides the provenance aspect of knowing exactly how a file was produced via metadata stored in the file itself, storing these options allows further automation of Analysis Productions (described in Section 3), as well as configuration consistency checks, as configuration options to run the DaVinci application on Spruced files can be deduced from the FSR. These configuration options include the detector geometry version and data-taking conditions, the file type, the input process (e.g., HLT2) and whether the file contains real or simulated data.
@@ 3.3 Production Testing and Submission
Should any monitored criteria exceed predefined thresholds, the tests will fail, and suggestions to fix the problem are provided to the user (via the merge request) on the AP GitLab platform. The productions, their status, and details of the validation are summarised on the platform, connected to aforementioned services, with curated job logs highlighting potential warnings and errors.
@@ 3.5 The apd Tool and Analysis Preservation
The apd tool provides a simple programming interface to the Analysis Productions system. It is a Python package published to standard repositories (such as PyPI [23] and conda-forge [24]) that allows the look-up of the physical file names of files created in the context of an Analysis Production, filtering them according to associated metadata (tags). Some of these tags are automatically created by the system (e.g., the data type), but analysts can also define custom tags that indicate the dataset properties or its intended use. Using apd, analysts can avoid keeping lists of physical file names that have time-limited validity and instead declare input data with information that represents its properties and intended use (e.g., the analysis name, the working group it is attached to, the data-taking year, the event type in case of simulated data, etc.).
@@
The ease of use is crucial to help with the adoption of any new system. The apd tool is installed in the default LHCb analysis environment which is available through lb-conda. The lb-conda tool defines LHCb Conda wrapper scripts [26] that provide access to Conda versions on CERN’s cvmfs [27]. The apd tool has few dependencies and can be installed on any machine independently of the other LHCb software applications if necessary. Using apd from Python scripts requires minimal effort from analysts, and also provides the functionality to cache metadata and files locally, to improve performance, or reproduce part of the analysis locally.
@@
Combining the use of apd to search for LHCb Analysis Production data files, lb-conda to ensure the reproducibility of the analysis environments and Snakemake to define the analysis workflows ensures comprehensive and reliable analysis preservation. The ease of use of this system has led to significant uptake from LHCb analysts.
@@ 4 The DaVinci Application
Analysis Productions uses the offline analysis software project DaVinci to create nTuples for further high-level analysis. The DaVinci application is built on the Gaudi [32] framework software package for processing High Energy Physics event data. In Run 3 DaVinci employs the purpose-built FunTuple component to facilitate the storage of event data in ROOT format, optimising it for subsequent offline analysis [33]. These nTuples may contain information about the selected signal candidates and the decay products such as particle kinematics, particle identification hypotheses, vertex fit qualities, etc. Within DaVinci one can run the DecayTreeFitter algorithm [34] on the signal candidates to constrain the particle momenta to point to particular production vertices and constrain the composite particle masses to improve the resolution of the reconstructed objects.
@@ 5 LHCb Offline Data Processing and Analysis in LHCb Upgrade II
  • Binning of variables for histogram based analyses.
Generally, these steps are trivially parallelisable and are well suited to grid productions. WLCG resources are experiment agnostic and can run any code/tool that can be packaged to cvmfs or deployed via a container. Portable analysis environments thus allow APs to run any application required for the above steps. In particular, the LHCb lb-conda tool that provides access to conda environments in cvmfs offers a comprehensive (versioned) default environment with all common HEP software packages, as well as the LHCb calibration tools. Alongside this, analysts can create customised environments specifically for their analysis, which are also versioned, enabling full analysis environment preservation.